# Cumulants and ordering of their ratios in 2D Potts models: Lessons for QCD?

### 19 Dec 2023 v1

@@ IV.1 Critical exponents
Table 2: The extracted values for position ($T_{C,L}$), height ($\chi_{2}(T_{C,L})$) and width ($\sigma_{C}$) of the peak of $\chi_{2}$ distribution as a function of temperature for $q$ = 3 state Potts model. The quoted values after ’$\pm$’ represent statistical uncertainties derived from the fit.
@@
We plot the higher-order susceptibilities, $\chi_{3}$ and $\chi_{4}$ for both $q=2$ and $q=3$ state Potts model in Fig. 3 as a function of temperature scaled by $T_{C,L}$. Since, these $\chi_{n}$’s are one order derivative of its preceding $\chi_{n-1}$ (for $n>2$), the sign of peak structure gets reversed in $\chi_{n}$ with respect to $\chi_{n-1}$. Similarly, $\chi_{5}$ and $\chi_{6}$ are shown as a function of $T/T_{C,L}$ for both the models in Fig. 4. The height of the peaks or dips in these higher-order susceptibilities are found to be lattice size dependent. They also follow a scaling as a function of $L$ as discussed in section II.2. In order to determine the scaling exponents of the higher-order susceptibilities, we fit the maximum or minimum value of $\chi_{n}$, depending on which is prominent, as a function of $L$ against the scaling relation: $\chi_{n}^{\text{max/min}}\propto L^{b}$. The scaling exponents of $\chi_{3}$, $\chi_{4}$, $\chi_{5}$ and $\chi_{6}$ obtained from the fit are referred as $e_{n}^{{}^{\prime\prime}}$. For $q=2$ state Potts model, the values of scaling exponents are given in Table 3, whereas for $q=3$ state Potts model, they are provided in Table 4. Additionally, we compare the scaling exponents (Eq. 14) obtained using exact values of $\gamma$ and $\nu$ as well as by using the estimates of $\gamma$ and $\nu$ (from scaling of $\chi_{2}$) obtained in this paper, and they are referred as $e_{n}$ and $e_{n}^{\prime}$ respectively in Tables 3 and 4. The numerically estimated scaling exponents, $e_{n}^{\prime\prime}$ for the higher-order susceptibilities show a good agreement with the analytically obtained values. We also note that determining the exponents from finite size scaling of $\chi_{n}$’s instead of using the numerical estimate of $\gamma$ and $\nu$ from $\chi_{2}$ result in smaller numerical errors.
@@
Table 3: Finite size scaling exponents of higher-order susceptibilities, $e_{n}^{{}^{\prime\prime}}$ obtained from fits to peak (or dip) values of $\chi_{n}$ as a function of lattice size in two-dimensional $q$ = 2 state Potts model. Here, $e_{n}$ is computed based on exact values of $\gamma$ and $\nu$ ($\gamma=7/4$ and $\nu=1$) in Eq. 14, while $e_{n}^{\prime}$ is derived using our estimates of $\gamma$ and $\nu$ discussed in Sec. IV.1. The quoted values after ’$\pm$’ represent statistical uncertainties derived from the fit.
@@
Table 4: Finite size scaling exponents of higher-order susceptibilities, $e_{n}^{{}^{\prime\prime}}$ obtained from fits to peak (or dip) values of $\chi_{n}$ as a function of lattice size in two-dimensional $q$ = 3 state Potts model. Here, $e_{n}$ is computed based on exact values of $\gamma$ and $\nu$ ($\gamma=13/9$ and $\nu=5/6$) in Eq. 14, while $e_{n}^{\prime}$ is derived using our estimates of $\gamma$ and $\nu$ discussed in Sec. IV.1. The quoted values after ’$\pm$’ represent statistical uncertainties derived from the fit.
Figure 5: Ratios of the susceptibilities, $\chi_{3}/\chi_{1}$, $\chi_{4}/\chi_{2}$, $\chi_{5}/\chi_{1}$ and $\chi_{6}/\chi_{2}$ for two-dimensional $q$ = 2 state Potts model with lattice sizes, L = 50 (left plot), 60 (middle plot), and 80 (right plot) are shown as a function of scaled temperature, $T/T_{C,L}$. The yellow shaded region represents the critical region, $T_{C,L}\pm\sigma_{C}$.
Figure 6: Ratios of the susceptibilities, $\chi_{3}/\chi_{1}$, $\chi_{4}/\chi_{2}$, $\chi_{5}/\chi_{1}$ and $\chi_{6}/\chi_{2}$ for two-dimensional $q$ = 3 state Potts model with lattice sizes, L = 50 (left plot), 60 (middle plot), and 80 (right plot) are shown as a function of scaled temperature, $T/T_{C,L}$. The yellow shaded region represents the critical region, $T_{C,L}\pm\sigma_{C}$.
IV.2 Ordering of Cumulant ratios
In this section, we delve into the discussion of whether the two-dimensional $q=2$ and $q=3$ state Potts models exhibit ordering in the susceptibility ratios, $\frac{\chi_{6}}{\chi_{2}}<\frac{\chi_{5}}{\chi_{1}}<\frac{\chi_{4}}{\chi_{2}}<\frac{\chi_{3}}{\chi_{1}}$. Figure 5 presents a comparison of susceptibility ratios, $\frac{\chi_{3}}{\chi_{1}}$, $\frac{\chi_{4}}{\chi_{2}}$, $\frac{\chi_{5}}{\chi_{1}}$, and $\frac{\chi_{6}}{\chi_{2}}$, in the $q$ = 2 state Potts model for lattice sizes of 50, 60, and 80 as a function of temperature scaled by $T_{C,L}$. The shaded yellow band corresponds to the critical region, $T_{C,L}\pm\sigma_{C}$, where the values of $\sigma_{C}$ are listed in Table 1. It is already clear from Figs. 3 and 4 that the scale of the susceptibilities differ by factors larger than $\sim\mathcal{O}(10^{3})$. Therefore, to investigate the inequalities among the susceptibility ratios (if it exists) thoroughly, the ratios are displayed pairwise in three panels (for easy inspection by eye). Furthermore, a meaningful discussion of the ordering of the ratio of susceptibilities is feasible in the critical region around $T_{C,L}$ as a result of the vast changes in the magnitudes of successively higher orders. A clear ordering, $\frac{\chi_{5}}{\chi_{1}}<\frac{\chi_{4}}{\chi_{2}}<\frac{\chi_{3}}{\chi_{1}}$ is observed for $T>T_{a}$ in the critical region from the top and middle panel plots of Fig. 5, where $T_{a}\sim 1.01T_{C,L}$ ($T_{a}\sim 1.005T_{C,L}$) for $L=50$ ($L=80$). But the ordering, $\frac{\chi_{6}}{\chi_{2}}<\frac{\chi_{5}}{\chi_{1}}$ is not observed for $T>T_{a}$. Instead, a reverse ordering is observed, with $\frac{\chi_{5}}{\chi_{1}}$ being smaller than $\frac{\chi_{6}}{\chi_{2}}$ for the entire $T>T_{C,L}$. Therefore, we find that the expected complete ordering as seen in [11] is not observed in any temperature range near the critical region in $q=2$ state Potts model. Only a truncated ordering of the susceptibility ratios up to $\frac{\chi_{5}}{\chi_{1}}$, i.e., $\frac{\chi_{5}}{\chi_{1}}<\frac{\chi_{4}}{\chi_{2}}<\frac{\chi_{3}}{\chi_{1}}$ holds true for $T>1.01T_{C,L}$ across all lattice sizes shown in Fig. 5. Such a truncated inequality is not observed for the entire temperature range, $T<T_{a}$.
@@ IV.2 Ordering of Cumulant ratios
A partial reversal of the ordering up to $\frac{\chi_{5}}{\chi_{1}}$, i.e., $\frac{\chi_{5}}{\chi_{1}}>\frac{\chi_{4}}{\chi_{2}}>\frac{\chi_{3}}{\chi_{1}}$ is seen in a narrow temperature range below the $T_{C,L}$, but a complete reversal is not observed for both $q=2$ and $q=3$ state Potts model. An important point to note is that regions of inequality exclusively lie in the critical region, which itself shrinks as in the infinite volume limit. This suggests that for infinite system size, we may not be able to establish any inequalities for the susceptibility ratios of magnetization in the spin models considered, and any observed orderings are a finite size effect. It would be interesting to check if any such volume dependence of the produced medium is seen in the relativistic heavy ion collisions.
V Conclusions
@@
LQCD predicts a specific hierarchy of net-baryon number susceptibility ratios near the quark-hadron transition temperature for small $\mu_{B}$: $\frac{\chi_{6}}{\chi_{2}}<\frac{\chi_{5}}{\chi_{1}}<\frac{\chi_{4}}{\chi_{2}}<\frac{\chi_{3}}{\chi_{1}}$. The STAR experiment tested these predictions for net-proton number cumulant ratios in Au+Au collisions at various center-of-mass energies. Cumulant ratios from 7.7 to 200 GeV exhibited the predicted hierarchy, while a reverse ordering was observed at 3 GeV. If these inequalities are universally applicable for any system undergoing phase transition, is what we tried to understand.
@@
In this study, we investigated higher-order susceptibilities of magnetization and their ratios in the critical regions of two-dimensional two and three-state Potts models. Utilizing the Wolff cluster algorithm to mitigate critical slowing down effects, simulations were conducted. We employed finite size scaling techniques to extract scaling exponents associated with higher-order susceptibilities, extending up to the sixth order. Additionally, we estimated these scaling exponents using the known critical exponents of the model, specifically the $\nu$ and $\gamma$ exponents. The obtained values from both methods exhibited a good agreement, indicating that the behavior of higher-order susceptibilities aligns well with our understanding of second-order phase transitions in the spin models.
@@
We observed an absence of a consistent order in susceptibility ratios both above and below the pseudo-critical point (pseudo-critical temperature). A truncated ordering, $\frac{\chi_{5}}{\chi_{1}}<\frac{\chi_{4}}{\chi_{2}}<\frac{\chi_{3}}{\chi_{1}}$ is observed in a small temperature range, greater than the pseudo-critical temperature for a given lattice size in both two and three-state Potts models. In a yet narrow temperature range just above the pseudo-critical temperature, the complete ordering hierarchy in the susceptibility ratios is noted in the three-state Potts models for $L=128$ (illustrated in Fig. 6). While a complete reversal of this ordering is not observed in either of the models, we acknowledge that a partial reversal of the ordering, $\frac{\chi_{5}}{\chi_{1}}>\frac{\chi_{4}}{\chi_{2}}>\frac{\chi_{3}}{\chi_{1}}$ is seen in a small temperature range below the pseudo-critical temperature. Furthermore, we found that the region suitable for making comparisons between these ratios diminishes notably as the lattice size (or “lattice volume”) increases, shrinking to a point in the infinite volume limit. Consequently, the orderings observed in the susceptibility ratios can be attributed to finite size effects in spin model.
The spin models considered here have only a second order phase transition. A more realistic simulation which will be in better analogy with the expectations for the QCD phase diagram entails employing $q=3$ state Potts model on a three-dimensional lattice and with a small but non-vanishing magnetic field. It will be interesting to see what changes does a first order phase transition or the cross over region beyond the critical magnetic field make to the above observations on the inequality structure.
Acknowledgements.
@@
We acknowledge the support of the Department of Atomic Energy (DAE) and the Department of Science and Technology (DST), Government of India. RVG expresses gratitude for the financial support from DAE through the Raja Ramanna Fellowship and the hospitality of the Department of Physics at IISER Bhopal. BM and SS acknowledge the financial support of DAE and DST. We also acknowledge the use of Garuda HPC facility at the School of Physical Sciences, NISER. SS would like to thank V K S Kashyap for technical help in running simulations, as well as Ashish Pandav and Debasish Mallick for insightful discussions regarding STAR results.

### 15 Apr 2026 v2

@@ 3 Numerical simulation
Table 2: The extracted values for position ($T_{C,L}$), height ($\chi_{2}(T_{C,L})$) and width ($\sigma_{C,L}$) of the peak of $\chi_{2}$ distribution on an $L^{2}$ lattice as a function of temperature for $q=3$ state Potts model. The quoted values after $\pm$’ represent statistical uncertainties derived from the fit.
In order to obtain independent spin configurations on the lattice, configurations were recorded at the interval of $2\tau$ where $\tau$ is the auto-correlation time that is obtained from the auto-correlation function (of $M$) defined in Eq. 16 using an exponential decay ansatz $\phi(t)\sim\mathrm{exp}\left({-\frac{t}{\tau}}\right)$.
4 Results
The top panels of Fig. 1 and Fig. 2 display the second-order susceptibility, denoted as $\chi_{2}$, for the $q=2$ and $q=3$ state Potts model respectively with varying lattice sizes, $L$. In both cases, $\chi_{2}$ exhibits a distinctive peak structure. As may be expected, the locations ($T_{C,L}$) and magnitudes ($\chi_{2}(T_{C,L})$) of these peaks vary with changes in the lattice size. Their scaling behavior can be utilized to extract the critical exponents, specifically $\gamma$ and $\nu$, which will be further elaborated upon in the subsequent section, 4.1.
Figure 1: (a) The second order susceptibility of magnetization for the 2d Ising model (equivalent to $q=2$ state Potts model) shown as a function of temperature. The results shown by different colored markers represent different lattice sizes ($L$). The dotted lines connecting the points are intended to guide the viewer’s eye. (b) The reduced temperature, $t^{\prime}=T(\chi_{2}^{\mathrm{max}})-T_{C}$ is plotted as a function of $1/L$, and fitted with the function $y=Ax^{1/\nu}$ to extract the value of $\nu$.
Figure 2: (a) The second order susceptibility of magnetization for the 2d $q=3$ state Potts model shown as a function of temperature. The results shown by different colored markers represent different lattice sizes ($L$). The dotted lines connecting the points are intended to guide the viewer’s eye. (b) The reduced temperature, $t^{\prime}=T(C_{2}^{\mathrm{max}})-T_{C}$ is plotted as a function of $1/L$, and fitted with the function $y=Ax^{1/\nu}$ to extract the value of $\nu$.
4.1 Critical exponents
The critical exponents, $\gamma$ and $\nu$ can be obtained using the scaling relations, $\chi_{2}(T_{C,L})\propto L^{-\gamma/\nu}$ and $T_{C,L}-T_{C}\propto L^{-1/\nu}$, where $T_{C,L}$ is the pseudo-critical temperature for a given lattice size, and $T_{C}$ represents the critical temperature of the model in the limit $L\to\infty$. The pseudo-critical temperature is an effective critical temperature observed in finite-sized systems, reflecting behavior similar to the critical point in infinite-sized systems. $T_{C,L}$ is defined as the location of the peak in the second order susceptibility, $\chi_{2}$ distribution as shown in Figs. 1 and  2. By fitting $\chi_{2}$ as a Gaussian function of the temperature, $T$, the peak location ($T_{C,L}$, $\chi_{2}(T_{C,L})$) is determined for each lattice size. The obtained values of $T_{C,L}$ and $\chi_{2}(T_{C,L})$ as a function of $L$ are tabulated in Tables 1 and 2 for $q=$ 2 and $q=$ 3 state Potts model respectively. Also, we define the critical region around $T_{C,L}$ for each lattice size by the width of the $\chi_{2}$ distribution peak, which is equal to the $\sigma$ of the fitted Gaussian distribution function, represented here as $\sigma_{C,L}$. Using the values of $T_{C,L}$, the difference $t^{\prime}=T_{C,L}-T_{C}$ is plotted as a function of $1/L$, and fitted with its scaling relations for both $q=2$ and $q=3$ state Potts model as shown in the bottom panels of Figs. 1 and 2 respectively. Alternatively, $T_{C}$ can be estimated by plotting directly $T_{C,L}$ against $L$ and fitting with $y=A+Bx^{-1/\nu}$. The resulting estimates ($T_{C}=2.2780\pm 0.0089$ for $q=2$ and $T_{C}=0.9953\pm 0.0001$ for $q=3$) are found to be consistent with the known values [19, 29] within uncertainties. For precise measurements of the critical exponents, however we employ the known values of $T_{C}$ in the former method and extracted the coefficient $\nu$. The peak position, $\chi_{2}(T_{C,L})$ as a function of $L$, is also fitted with its scaling relation. Therefore, the extracted values of $\gamma$ and $\nu$ from the fits with the scaling relations are as follows: for $q=2$, $\gamma=1.758\pm 0.060$, $\nu=1.019\pm 0.034$ whereas for $q=3$, $\gamma=1.435\pm 0.014$, $\nu=0.830\pm 0.008$. The excellent agreement observed between the calculated values of $\gamma$ and $\nu$ and their theoretical counterparts affirms the validation of our simulation.
@@ 4.1 Critical exponents
We plot the higher-order susceptibilities, $\chi_{3}$ and $\chi_{4}$ for both $q=2$ and $q=3$ state Potts model in Fig. 3 as a function of temperature scaled by $T_{C,L}$. Since, these $\chi_{n}$’s are one order derivative of its preceding $\chi_{n-1}$ (for $n>2$), the sign of peak structure gets reversed in $\chi_{n}$ with respect to $\chi_{n-1}$. Similarly, $\chi_{5}$ and $\chi_{6}$ are shown as a function of $T/T_{C,L}$ for both the models in Fig. 4. The height of the peaks or dips in these higher-order susceptibilities are found to be lattice size dependent. They also follow a scaling as a function of $L$ as discussed in section 2.2. In order to determine the scaling exponents of the higher-order susceptibilities, we fit the maximum or minimum value of $\chi_{n}$, depending on which is prominent, as a function of $L$ against the scaling relation: $\chi_{n}^{\mathrm{max/min}}\propto L^{b}$. The scaling exponents of $\chi_{3}$, $\chi_{4}$, $\chi_{5}$ and $\chi_{6}$ obtained from the fit are referred as $e_{n}^{{}^{\prime\prime}}$. For $q=2$ state Potts model, the values of scaling exponents are given in Table 3, whereas for $q=3$ state Potts model, they are provided in Table 4. Additionally, we compare the scaling exponents (Eq. 15) obtained using exact values of $\gamma$ and $\nu$ as well as by using the estimates of $\gamma$ and $\nu$ (from scaling of $\chi_{2}$) obtained in this paper, and they are referred as $e_{n}$ and $e_{n}^{\prime}$ respectively in Tables 3 and 4. The numerically estimated scaling exponents, $e_{n}^{\prime\prime}$ for the higher-order susceptibilities show a good agreement with the analytically obtained values. We also note that determining the exponents from finite size scaling of $\chi_{n}$’s instead of using the numerical estimate of $\gamma$ and $\nu$ from $\chi_{2}$ result in smaller numerical errors.
Figure 3: The third and fourth order susceptibilities of magnetization, $\chi_{3}$ and $\chi_{4}$, are shown for different lattice sizes as a function of temperature scaled by critical temperature of the corresponding lattice, $T/T_{C,L}$. The left side plot belongs to 2d $q=2$ state Potts model whereas the right side plot is for $q=3$ state Potts model. The upper panel represents $\chi_{3}$ and the bottom panel represents $\chi_{4}$ [where $\chi_{3}$ and $\chi_{4}$ divided by a factor $10^{5}$ ($10^{6}$) and $10^{8}$ ($10^{9}$) respectively for $q=2$ ($q=3$)].
Figure 4: The fifth and sixth order susceptibilities of magnetization, $\chi_{5}$ and $\chi_{6}$, are shown for different lattice sizes as a function of temperature scaled by critical temperature of the corresponding lattice, $T/T_{C,L}$. The left side plot belongs to 2d $q=2$ state Potts model whereas the right side plot is for $q=3$ state Potts model. The upper panel represents $\chi_{5}$ and the bottom panel represents $\chi_{6}$ [where $\chi_{5}$ and $\chi_{6}$ divided by a factor $10^{12}$ ($10^{13}$) and $10^{16}$ ($10^{17}$) respectively for $q=2$ ($q=3$)].
@@
Table 3: Finite size scaling exponents of higher-order susceptibilities, $e_{n}^{{}^{\prime\prime}}$ obtained from fits to peak (or dip) values of $\chi_{n}$ as a function of lattice size in 2d $q=2$ state Potts model. Here, $e_{n}$ is computed based on exact values of $\gamma$ and $\nu$ ($\gamma=7/4$ and $\nu=1$) in Eq. 15, while $e_{n}^{\prime}$ is derived using our estimates of $\gamma$ and $\nu$ discussed in Sec. 4.1. The quoted values after $\pm$’ represent statistical uncertainties derived from the fit.
@@
Table 4: Finite size scaling exponents of higher-order susceptibilities, $e_{n}^{{}^{\prime\prime}}$ obtained from fits to peak (or dip) values of $\chi_{n}$ as a function of lattice size in 2d $q=3$ state Potts model. Here, $e_{n}$ is computed based on exact values of $\gamma$ and $\nu$ ($\gamma=13/9$ and $\nu=5/6$) in Eq. 15, while $e_{n}^{\prime}$ is derived using our estimates of $\gamma$ and $\nu$ discussed in Sec. 4.1. The quoted values after $\pm$’ represent statistical uncertainties derived from the fit.
4.2 Ordering of Cumulant ratios
In this section, we delve into the discussion of whether the 2d $q=2$ and $q=3$ state Potts models exhibit ordering in the susceptibility ratios, $\frac{\chi_{6}}{\chi_{2}}<\frac{\chi_{5}}{\chi_{1}}<\frac{\chi_{4}}{\chi_{2}}<\frac{\chi_{3}}{\chi_{1}}$. While it may apparently seem reasonable to only compare susceptibility ratios that are proportional to the same powers of $L$ (such as $\frac{\chi_{6}}{\chi_{2}}$ and $\frac{\chi_{5}}{\chi_{1}}$ proportional to $L^{4}$, or $\frac{\chi_{4}}{\chi_{2}}$ and $\frac{\chi_{3}}{\chi_{1}}$ proportional to $L^{2}$), we cannot overlook the terms $C_{2}^{n}$ and $\frac{\partial^{n}Y}{\partial h_{r}^{n}}\bigg|_{h_{r}=0}$ in Eq. 13. In Fig.5, we plot the double ratio of susceptibilities, $\frac{\chi_{4}}{\chi_{2}}/\frac{\chi_{3}}{\chi_{1}}$ and $\frac{\chi_{6}}{\chi_{2}}/\frac{\chi_{5}}{\chi_{1}}$, for the $q=2$ state Potts model as a function of temperature scaled by $T_{C,L}$ in the top and bottom panels respectively. Similar results for the $q=3$ state Potts model can be seen in Fig.6. In both the $q=2$ and $q=3$ state Potts models, $\chi_{6}/\chi_{2}$ ($\chi_{4}/\chi_{2}$) is not always less than $\chi_{5}/\chi_{1}$ ($\chi_{3}/\chi_{1}$) across the studied temperature range. Despite the cancellation of $L$ dependence in the double ratios, the influence of $C_{2}^{n}$ and $\frac{\partial^{n}Y}{\partial h_{r}^{n}}\bigg|_{h_{r}=0}$ is significant in determining the inequality among the ratios. Therefore, it is crucial to have a comprehensive comparison of the entire set of inequalities in these spin models.
Figure 5: Double ratios $(\chi_{6}/\chi_{2})/(\chi_{5}/\chi_{1})$ and $(\chi_{4}/\chi_{2})/(\chi_{3}/\chi_{1})$ as a function of temperature scaled by critical temperature of the corresponding lattice, $T/T_{C,L}$ in $q=2$ state Potts model.
Figure 6: Double ratios $(\chi_{6}/\chi_{2})/(\chi_{5}/\chi_{1})$ and $(\chi_{4}/\chi_{2})/(\chi_{3}/\chi_{1})$ as a function of temperature scaled by critical temperature of the corresponding lattice, $T/T_{C,L}$ in $q=3$ state Potts model.
@@ 4.2 Ordering of Cumulant ratios
Figure 7: The difference among ratios of the susceptibilities, ($\chi_{3}/\chi_{1}-\chi_{4}/\chi_{2}$), ($\chi_{4}/\chi_{2}-\chi_{5}/\chi_{1}$) and ($\chi_{5}/\chi_{1}-\chi_{6}/\chi_{2}$) for 2d $q=2$ state Potts model with lattice sizes, L = 50 (left plot), 60 (middle plot), and 80 (right plot) are shown as a function of scaled temperature, $T/T_{C,L}$. The yellow shaded band indicates the critical region defined by $T_{C,L}\pm 2\sigma_{C,L}$, while the grey hatched region highlights the narrower interval $T_{C,L}\pm\sigma_{C,L}$.
These differences exhibit clear sign changes in the vicinity of $T_{C,L}$. A clear ordering, $\frac{\chi_{5}}{\chi_{1}}<\frac{\chi_{4}}{\chi_{2}}<\frac{\chi_{3}}{\chi_{1}}$, is observed on the high-temperature side ($T>T_{C,L}$), as indicated by the positive values of the corresponding differences in the top and middle panels of Fig. 7, consistently across all lattice sizes. This ordering sets in at $T\gtrsim T_{a}$, with $T_{a}\approx 1.01T_{C,L}$, and is observed within both the $1\sigma$ and $2\sigma$ regions. In contrast, the expected hierarchy $\frac{\chi_{6}}{\chi_{2}}<\frac{\chi_{5}}{\chi_{1}}$ is not observed across most of the critical region, where $\frac{\chi_{5}}{\chi_{1}}-\frac{\chi_{6}}{\chi_{2}}<0$. Only near the upper edge of the $2\sigma$ band on the high-temperature side does this difference become positive, thereby satisfying the complete ordering. For $T<T_{C,L}$, no consistent ordering is observed within either the $1\sigma$ or $2\sigma$ regions, although a reversed ordering, $\frac{\chi_{5}}{\chi_{1}}>\frac{\chi_{4}}{\chi_{2}}>\frac{\chi_{3}}{\chi_{1}}$, appears over a temperature interval near the lower edge of the $1\sigma$ band. These results demonstrate that the hierarchy among susceptibility ratios is not robust within the critical region and is only partially realized over a limited temperature interval above $T_{C,L}$. This lack of a robust, extended hierarchy in the $q=2$ state Potts model suggests that such cumulant-ratio inequalities may not be generic consequences of continuous phase transitions in finite systems.
Figure 8: The difference among ratios of the susceptibilities, ($\chi_{3}/\chi_{1}-\chi_{4}/\chi_{2}$), ($\chi_{4}/\chi_{2}-\chi_{5}/\chi_{1}$) and ($\chi_{5}/\chi_{1}-\chi_{6}/\chi_{2}$) for 2d $q=3$ state Potts model with lattice sizes, L = 48 (left plot), 80 (middle plot), and 128 (right plot) are shown as a function of scaled temperature, $T/T_{C,L}$. The yellow shaded band indicates the critical region defined by $T_{C,L}\pm 2\sigma_{C,L}$, while the grey hatched region highlights the narrower interval $T_{C,L}\pm\sigma_{C,L}$.
[…]
Taken together, the behavior observed for both $q=2$ and $q=3$ models indicates that a partial reversal of the ordering up to $\frac{\chi_{5}}{\chi_{1}}$, i.e., $\frac{\chi_{5}}{\chi_{1}}>\frac{\chi_{4}}{\chi_{2}}>\frac{\chi_{3}}{\chi_{1}}$ occurs below $T_{C,L}$, while a complete reversal remains absent in both cases. An important point to note is that regions of inequality exclusively lie in the critical region, which itself shrinks in the infinite volume limit. This suggests that for infinite system size, we may not be able to establish any inequalities for the susceptibility ratios of magnetization in the spin models considered, and any observed orderings are a finite size effect. However, since the studied inequalities are relevant to LQCD and experimental heavy-ion collisions, both of which deal with systems with finite volume (moderate $L/\xi$), this work holds significant importance. Our preliminary determinations of the correlation lengths suggest that the ratio of $L$ to the correlation length ($\xi$), denoted as $L/\xi$, ranges between approximately 9.1 and 13.3 for the lattice systems studied in this work. Notably, systems associated with LQCD studies and experimental heavy-ion collisions exhibit $L/\xi$ values lower than those obtained in our study, indicating that finite size effects also play a crucial role in shaping the observed behavior in those systems. It would be intriguing to explore whether the volume dependency observed for the inequalities in spin models also holds true for the produced medium in relativistic heavy-ion collisions.
5 Conclusions
@@
A specific hierarchy of net-baryon number susceptibility ratios near the quark-hadron transition temperature at small $\mu_{\mathrm{B}}$, namely $\frac{\chi_{6}}{\chi_{2}}<\frac{\chi_{5}}{\chi_{1}}<\frac{\chi_{4}}{\chi_{2}}<\frac{\chi_{3}}{\chi_{1}}$, has been proposed in the literature [5, 6, 14]. The STAR experiment tested these predictions for net-proton number cumulant ratios in Au+Au collisions at various center-of-mass energies. Cumulant ratios from 7.7 to 200 GeV exhibited the predicted hierarchy, while a reverse ordering was observed at 3 GeV. If these inequalities are universally applicable for any system undergoing phase transition, is what we tried to understand.
@@
Although the QCD critical point belongs to the three-dimensional Ising universality class, the qualitative behavior of fluctuations near criticality and their sensitivity to finite system size can be explored in simpler statistical systems such as spin models. In this study, we investigated higher-order susceptibilities of magnetization and their ratios in the critical regions of 2d two and three-state Potts models. Utilizing the Wolff cluster algorithm to mitigate critical slowing down effects, simulations were conducted. We employed finite size scaling techniques to extract scaling exponents associated with higher-order susceptibilities, extending up to the sixth order. Additionally, we estimated these scaling exponents using the known critical exponents of the model, specifically the $\nu$ and $\gamma$ exponents. The obtained values from both methods exhibited a good agreement, indicating that the behavior of higher-order susceptibilities aligns well with our understanding of second-order phase transitions in the spin models.
We do not observe a consistent ordering of susceptibility ratios either above or below the pseudo-critical temperature. A truncated hierarchy, $\frac{\chi_{5}}{\chi_{1}}<\frac{\chi_{4}}{\chi_{2}}<\frac{\chi_{3}}{\chi_{1}}$, appears only on the high-temperature side within a narrow temperature interval in the $1\sigma$ band of the critical region for both $q=2$ and $q=3$ state Potts models. The hierarchy involving $\frac{\chi_{6}}{\chi_{2}}$ is not satisfied within the $1\sigma$ region. A complete ordering of the susceptibility ratios is realized for both models only in an even narrower temperature interval on the high-temperature side outside the $1\sigma$ region, near the upper edge of the $2\sigma$ band; this interval shrinks with increasing lattice size. The statistical robustness of this feature is assessed using resampling-based uncertainty estimates. However, given the limited width of the temperature interval in which this ordering is observed, it should be interpreted as a finite-size effect whose visibility depends sensitively on statistical precision and analysis details.
@@
Below the pseudo-critical temperature, a complete reversal of this ordering is not observed in either model; only a partial reversal, $\frac{\chi_{5}}{\chi_{1}}>\frac{\chi_{4}}{\chi_{2}}>\frac{\chi_{3}}{\chi_{1}}$, is seen within a small temperature range. Furthermore, the temperature region over which meaningful comparisons between susceptibility ratios can be made diminishes notably as the lattice size (or “lattice volume”) increases, and is expected to shrink to a point in the infinite volume limit. These results therefore demonstrate that while critical scaling is a universal feature of finite systems near second-order phase transitions, the ordering of higher-order cumulant ratios is not, highlighting the limitations of universality arguments for such observables. Instead, such ordering emerges only within narrow temperature intervals and weakens with increasing system size, indicating a strong sensitivity to model details and finite-size effects. Consequently, the orderings observed in the susceptibility ratios can be attributed to finite size effects in spin models.
The 2d spin models considered here have only a second order phase transition. A more realistic analog to the QCD phase diagram is provided by the 3d $q=3$ state Potts model with a small, non-vanishing magnetic field $h$, where the first-order transition at $h=0$ weakens into a critical endpoint as $h\to h_{c}$ (critical magnetic field). This establishes a qualitative correspondence between the $(T,h)$ plane—encompassing first-order, critical, and crossover regions—and the QCD $(T,\mu_{\mathrm{B}})$ plane. In such a framework, the behavior of higher-order cumulants may depend on the trajectory along which they are evaluated. For instance, following a path at fixed temperature $T_{fc}$ on the first-order line by varying only $h$ does not necessarily pass through or probe the vicinity of the critical point, which may thereby affect the ordering of cumulant ratios. This suggests that the absence of a robust hierarchy in the present study may be influenced by the specific path (fixed $h=0$) in parameter space, indicating that the ordering of cumulant ratios may not be a universal feature even within a given model. It would therefore be of interest to investigate how the hierarchy of cumulants evolves across the different regions of the 3d Potts $(T,h)$ plane. In addition, for a meaningful comparison with experimental measurements in heavy-ion collisions, global conservation laws must be taken into account. Baryon number conservation, which holds event-by-event, can modify higher-order net-baryon cumulants in a non-linear manner. Also, since measurements in experiment are performed within a limited kinematic acceptance, the observed cumulant ratios depend on the fraction of the system that is probed, and the resulting hierarchy may be altered. Such effects are not included in the present spin-model study and should be considered for a more direct comparison with experimental data.
@@
We acknowledge the support of the Department of Atomic Energy (DAE) and the Department of Science and Technology (DST), Government of India. RVG expresses gratitude for the financial support from DAE through the Raja Ramanna Fellowship and the hospitality of the Department of Physics at IISER Bhopal. BM and $S$ acknowledge the financial support of DAE and DST. We also acknowledge the use of Garuda HPC facility at the School of Physical Sciences, NISER. $S$ would like to thank V K S Kashyap for technical help in running simulations, as well as Ashish Pandav and Debasish Mallick for insightful discussions regarding STAR results.
6 Data availability statement
Our data that underlie the findings are presented graphically within the article and are available in tabular format in the Supplementary Material.
Appendix A Ratio of cumulants
Figure 9: Ratios of the susceptibilities, $\chi_{3}/\chi_{1}$, $\chi_{4}/\chi_{2}$, $\chi_{5}/\chi_{1}$ and $\chi_{6}/\chi_{2}$ for 2d $q=2$ state Potts model with lattice sizes, L = 50 (left plot), 60 (middle plot), and 80 (right plot) are shown as a function of scaled temperature, $T/T_{C,L}$. The yellow shaded band indicates the critical region defined by $T_{C,L}\pm 2\sigma_{C,L}$, while the grey hatched region highlights the narrower interval $T_{C,L}\pm\sigma_{C,L}$.
Figure 10: Ratios of the susceptibilities, $\chi_{3}/\chi_{1}$, $\chi_{4}/\chi_{2}$, $\chi_{5}/\chi_{1}$ and $\chi_{6}/\chi_{2}$ for 2d $q=3$ state Potts model with lattice sizes, L = 48 (left plot), 80 (middle plot), and 128 (right plot) are shown as a function of scaled temperature, $T/T_{C,L}$. The yellow shaded band indicates the critical region defined by $T_{C,L}\pm 2\sigma_{C,L}$, while the grey hatched region highlights the narrower interval $T_{C,L}\pm\sigma_{C,L}$.